OpenStudy (toxicsugar22):

how do u solve 4x+5y=7 7y-x=23

OpenStudy (anonymous):

u can solve it by substitution method or by elimination method or by cross-multiplication method.... choice is yours ☺

OpenStudy (toxicsugar22):

well can u show me any chice lets say elimatnion

OpenStudy (anonymous):

multiply bottom equation by 4 you will do this so that it you can eliminate the x variable when adding both equations so you will be left with only one equation on the y variable 4x-4x+5y+28y=7+92 and that is 33y=99 and you should be able to solve that one and then replace the result on any of the other original equations to get the value for x

OpenStudy (anonymous):

4x + 5y = 7 7y - x = 23 Take the first equation and solve for x in terms of y: 4x + 5y = 7 4x = 7 - 5y x = (7 - 5y)/4 Now plug this value of x into the other equation: 7y - x = 23 7y - ((7 - 5y)/4) = 23 Solve for y: 7y * (4/4) - ((7 - 5y)/4) = 23 28y/4 - (7 - 5y)/4 = 23 (28y - 7 + 5y)/4 = 23 28y - 7 + 5y = 92 33y - 7 = 92 33y = 99 y = 3 plug y into one of the equations to solve for x: 4x + 5y = 7 4x + 5(3) = 7 4x = -8 x = -2

OpenStudy (dumbcow):

@harkirat what is the cross-multiplication method?

OpenStudy (anonymous):

wait i'll come back to reply....

OpenStudy (anonymous):

if u have a system of linear equations in two variables x and y as follows: a1x + b1y + c1 = 0 and a2x + b2y + c2 = 0 then x y 1 ------------ = -------------- = -------------- b1c2-b2c1 c1a2 - c2a1 a1b2 - a2b1

OpenStudy (dumbcow):

ok thanks, i see how that works just never seen that method used before

OpenStudy (toxicsugar22):

thnaks

OpenStudy (dumbcow):

its basically cramers rule when using matrices

OpenStudy (dumbcow):

thats usually too advanced for h.s algebra students

OpenStudy (anonymous):

to remember this easily we write the coefficients and variables as shown in the attached file (and you will see why we call it cross-multiplication method

OpenStudy (anonymous):

In India, I never heard of Cramer's rule !!! We just teach it as cross-multiplication method.....

OpenStudy (dumbcow):

o ok...yeah sorry US here obviously education curriculum is different

OpenStudy (anonymous):

yeah some "cultural" differences.... I see you people using many shortcuts and calculators etc for many question, which we never do in India...

OpenStudy (anonymous):

In school we have to do everything with full/proper method. However, in competitive exams, u can use shortcuts for calculations ☺

OpenStudy (anonymous):

4x + 5y = 7 7y - x = 23 Take the first equation and solve for x in terms of y: 4x + 5y = 7 4x = 7 - 5y x = (7 - 5y)/4 Now plug this value of x into the other equation: 7y - x = 23 7y - ((7 - 5y)/4) = 23 Solve for y: 7y * (4/4) - ((7 - 5y)/4) = 23 28y/4 - (7 - 5y)/4 = 23 (28y - 7 + 5y)/4 = 23 28y - 7 + 5y = 92 33y - 7 = 92 33y = 99 y = 3 plug y into one of the equations to solve for x: 4x + 5y = 7 4x + 5(3) = 7 4x = -8 x = -2

OpenStudy (dumbcow):

ideally we aim to do the same, however, there are many standards that have to taught and there are now much more standardized tests for students so teachers just give them the shortcuts

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Cramers_rule

OpenStudy (anonymous):

Well, I find the Indian schooling system being modified and I do not like it at all. In my time if one scored 70% marks, one was considered a brilliant student. At that time the all-India topper used to score about 85% marks. Nowadays students r scoring 90% plus marks and not getting admission in colleges because the percentage of students scoring 90% plus marks has increased like plague.....

OpenStudy (anonymous):

thanks for the link.. :))